OpenStudy (anonymous):

Given f(x)=2x+5 and g(x)=8x+2 then what is (f−g)(−9)) ?

OpenStudy (anonymous):

@UsukiDoll think you could help please ?

Vocaloid (vocaloid):

first off, (f-g)(x) = f(x)-g(x) = __ ?

OpenStudy (anonymous):

8x+2

Vocaloid (vocaloid):

not quite f(x) = 2x + 5 g(x) = 8x + 2 f(x) - g(x) = ?

OpenStudy (anonymous):

-6x+7?

Vocaloid (vocaloid):

not quite, check your math one more time f(x) - g(x) = (2x + 5) - (8x + 2) = ?

OpenStudy (anonymous):

-6x+3?

Vocaloid (vocaloid):

good! so we know that (f-g)(x) = -6x + 3 so what is (f-g)(-9)= ?

OpenStudy (anonymous):

2x+14?

Vocaloid (vocaloid):

uh... not quite x = -9 (f-g)(x) = -6x + 3 = ?

OpenStudy (anonymous):

8x+11?

Vocaloid (vocaloid):

x = -9 means replace "x" with "-9" whenever you see it (f-g)(x) = -6x + 3 = ?

Vocaloid (vocaloid):

if it's still unclear... (f-g)(x) = -6x + 3 = -6*(-9) + 3 = ?

OpenStudy (anonymous):

57?

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

so thats the answer?

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

could you help me with a few more please ? @Vocaloid

Vocaloid (vocaloid):

I need to get going, let me tag @Nnesha

OpenStudy (anonymous):

@Nnesha think you could help me with this one please Match the correct graph to the function shown. f(x)=2sqrt3-4x +3

OpenStudy (usukidoll):

@Vocaloid thanks for helping. I'm super exhausted right now and just want to browse OpenStudy.